Can someone help solve this?
Let H(t)=4+4sin(pi t/2). Approximate to the nearest hundredth
all values t in the interval [0,4 pi] for which Ht=6.

Answer by venugopalramana(3286)

Can someone help solve this?
Let H(t)=4+4sin(pi t/2). Approximate to the nearest hundredth
all values t in the interval [0,4 pi] for which Ht=6.
6=4+4SIN(PI*T/2)
SIN(PI*T/2)=(6-4)/4=1/2
PI*T/2=PI/6....OR.....PI-PI/6....OR....2PI+PI/6...........
T=1/3.....OR....5/3.........13/3.
IN THE RANGE OF 0 TO 4PI
T=1/3....OR....5/3....